DIRECT POTENTIAL FITTING FOR THE A^{3}Π_{1u} and X^{1}Σ_{g}^{+} STATES OF Br_{2}

\underline{TOKIO YUKIYA}, NOBUO NISHIMIYA, MASAO SUZUKI, Department of Electronics and Information Technology, Tokyo Polytechnic University, Iiyama 1583, Atsugi City, Kanagawa 243-0297, Japan; ROBERT J. LE ROY, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

Bromine dimer has been studied by many researchers in various wavelength regions. The A^{3}Π_{1u} and X^{1}Σ_{g}^{+} states have been well observed using magnetic rotation spectroscopy,^{a} by laser-induced fluorescence,^{b}^{,}^{c} by laser absorption,^{d} by Fourier transform absorption,^{e} and by UV emission.^{f} This yields a data set consisting of 16916 transitions in which the observed vibrational levels for the X^{1}Σ_{g}^{+} and A^{3}Π_{1u} states span 83% and 99% of the potential well depths, respectively, with the highest observed vibrational level of the A^{3}Π_{1u} state lying only 2 cm^{−1} below the dissociation limit.^{a} In order to provide the most compact and comprehensive description of these data, and the ability to make reliable predictions outside their range, we have chosen to perform a “direct potential fit” (DPF), rather than a conventional Dunhamexpansion analysis. In particular, accurate analytic potential energy functions for the A^{3}Π_{1u} and X^{1}Σ_{g}^{+} states are determined from a combined-isotopologue DPF analysis that also yields the electronic isotope shift, the Ω-doubling radial strength function, and an experimental value for the long-range inverse-power C_{5} constant of the A^{3}Π_{1u} state, as well as centrifugal Born-Oppenheimer Breakdown (BOB) functions for both states. To reveal characteristics of the A^{3}Π_{1u} state, band constants calculated from these potentials are compared with those determined from a conventional parameter-fitting analysis reported by Coxon.^{g}

a C. D. Boone, PhD Thesis, University of British Columbia (1999).

b C. Focsa et al., J. Mol. Spectrosc. 200, 104 (2000).

c D. J. Postell et al., to be published.

d N. Nishimiya el al., Columbus Meeting, paper WH02 (2005).

e S. Gerstenkorn et al., J. Physique, 48, 1685 (1987).

f P. Venkateswarlu et al., J. Mol. Spectrosc. 96, 247 (1982).

g J. A. Coxon, J. Mol. Spectrosc. 41, 548 (1972).